OpenStudy (anonymous):

Calculate the standard deviation of the data below: 8, 24, 6, 15 Question 1 options: a) 7.05 b) 7.25 c) 8.04 d) 8.14

OpenStudy (mertsj):

Have you calculated the mean?

OpenStudy (mertsj):

That is your first step.

OpenStudy (isaiah.feynman):

Calculate the mean first, then the variance. After getting the variance, take its square root and you get the standard deviation.

OpenStudy (isaiah.feynman):

\[Mean =\frac{\Sigma x }{ n } \] \[Variance = \frac{ \Sigma(x-mean) }{ n-1 } \] \[SD=\sqrt{variance}\]

OpenStudy (isaiah.feynman):

x is the individual numbers. In statistics they are called observations.

OpenStudy (anonymous):

the mean is 53

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (isaiah.feynman):

Mean is sum of the numbers divided by the number of numbers. So 53/4. There are four numbers there.

OpenStudy (anonymous):

that is 13.25

OpenStudy (isaiah.feynman):

Thats the mean!

OpenStudy (anonymous):

okay how u find the variance

OpenStudy (isaiah.feynman):

Use the formula I showed!

OpenStudy (amistre64):

subtract the mean from each data point, square the results and take the mean of the squares.

OpenStudy (amistre64):

8, 24, 6, 15 8-m, 24-m, 6-m, 15-m (8-m)^2, (24-m)^2, (6-m)^2, (15-m)^2 (8-m)^2 + (24-m)^2 + (6-m)^2 + (15-m)^2 = K if the dataset is a population, var = K/4 if its a sample, var = K/(4-1) and standard deviation is just the sqrt of variance